1.00 mL of a 3.50  10–4 M solution of oleic acid is diluted with 9.00 mL of petroleum ether,

forming solution A. Then 2.00 mL of solution A is diluted with 8.00 mL of petroleum ether,

forming solution B. What is the concentration of solution B?


1
Expert's answer
2021-10-04T05:09:56-0400

Total volume of solution A = (9.00+1.00) mL = 10.00 mL


According to law of dilution,


where C1 and C2are initial and final concentration respectively. V1 and V2 are initial and final volume respectively.

C1=3.55×10-4M

V1 = 1.00 mL, V2= 10.00 mL








So, concentration of solution A =




Solution B: Total volume of solution B = (2.00+8.00) mL = 10.00 mL

Similarly as above,





Therefore the concentration of solution B= 7.10×10-6M